when trying to find the percent of an number where do you start. example:
2% out of 100
3%of 100
6% of 65

Answer by josmiceli(19441)

per cent means "per one hundred". Cent is Latin for 100.
So, when you see "%" think of "/100"
2% = 2/100 or in decimals, .02
You are told to take 2% of a number, and that number happens to be 100.
So multiply .02 by 100
.02 * 100 = 2
likewise,
.03 * 100 = 3
Then you're asked for 6% of 65
.06 * 65 = 3.9